what is the answer to 3/(v+1) + 6/(v-7) equal? i need to know how to show work too.
Answer by tommyt3rd(5050)

3/(v+1) + 6/(v-7) =
(3(v-7)+6(v+1))/((v-7)(v+1))=
(9v-15)/((v-7)(v+1))